LaborTechs, LLC supplies contract labor (certified mechanics and technicians) to aircraft maintenance, repair, overhaul, and manufacturing companies. For the most recent year, LaborTechs budgeted to supply 86,000 hours of contract of labor. Its variable costs are $35 per hour and its fixed costs are $430,000. Roger Newton, the general manager, has always used a cost-plus approach for pricing contract labor at full cost plus 20%.

Required:

  1. Calculate the price per hour that LaborTechs will charge using Newton’s cost-plus approach to pricing.
  2. The marketing manager at LaborTechs has studied the relation between price and demand for LaborTechs' services and supplies the following schedule of demand at different prices:

Price per hour

Estimated Demand (hours)

$44.00

127,000

$46.00

107,000

$48.00

86,000

$50.00

73,000

$52.00

60,000

LaborTechs has capacity to meet any of these demand levels without incurring any additional fixed costs. Using this additional information, calculate the price per hour that LaborTechs should charge to maximize net operating income.

  1. Write a succinct statement (at least one sentence but no more than one paragraph) comparing the advantages of the cost-plus pricing approach in question 1 and the pricing approach in question 2.
Answer #1

1. Calculation of price per hours to be charged.

a. Total budgted hours of supply = 86,000

b. Variable cost per hour = 35

c. Total fixed cost = 430,000

d. Allocated fixed cost per hour = 430,000/86000 = 5

e. Total cost per hour = 35+ 5 = 40

Profit to be charged = 20%

Price to be charged = cost + profit (20% of cost) = 40+ (20%*40) = 42

Cost to be charged as per cost plus method = 42

2. Please refer the table attached.

Maximum operating profit at $46.

3. Cost plus pricing approach ensure that a fixed profit is earned on each transaction and is simple to calculate.

Whereas in demand based pricing approach business earns the maximum profit based on its demand in the market.